the average grade of a group of 24 alegbra students is 85%, and the standard deviation is known to be 4%, find the 95% confidence interval for the true mean of the grades
Answer #1

n = 24     

x-bar = 85     

s = 4     

% = 95     

Standard Error, SE = s/√n =    4/√24 = 0.816496581

Degrees of freedom = n - 1 =   24 -1 = 23   

t- score = 2.068657599     

Width of the confidence interval = t * SE =     2.06865759861054 * 0.816496580927726 = 1.689051856

Lower Limit of the confidence interval = x-bar - width =      85 - 1.68905185637567 = 83.31094814

Upper Limit of the confidence interval = x-bar + width =      85 + 1.68905185637567 = 86.68905186

The confidence interval is [83.31%, 86.69%]